$SPAD/src/input negfloats.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool negfloats.output
)set message test on
)set message auto off
)clear all
--S 1 of 3
truncate(-9.6571)
--R
--R
--R   (1)  - 9.0
--R                                                                  Type: Float
--E 1
--S 2 of 3
fractionPart(-3.432)
--R
--R
--R   (2)  - 0.432
--R                                                                  Type: Float
--E 2
--S 3 of 3
round(-9.6571)
--R
--R
--R   (3)  - 10.0
--R                                                                  Type: Float
--E 3
)spool
)lisp (bye)

———-

References^{[1] nothing}

4